1. Find the following:

GIVEN:

Density of Water: 0.997770 g/cm3
Mass of Water: 53.959 g
Mass of Salt (NaCl) Solution: 66.114g

a.) Density of the salt solution:

b.) Molarity of the salt in the solution
Answer #1

Volume = mass/density

Volume of water =53.959/ 0.99777= 54.08 cm^3

The salt solution will have the same volume ( as volume of salt is negligible).

a) density of salt solution= mass of salt solution/ volume of salt solution

= 66.114/54.08 = 1.222 g/cm^3

b) mass of salt = mass of salt solution - mass of water = 66.114-53.959 = 12.155 g

Molar mass of NaCl = 58.44 g/mol

Moles of NaCl = mass/ molar mass = 12.155/58.44 = 0.208 mol

Molarity = moles of solute / volume of solution in L = 0.208/0.054 = 3.85 M